If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200


Usethemultiplierapplicabletoexport,toexplainhowa100–billiondeclineindemand

forexportcouldaffecttheeconomy’s:

(i)

Balanceofpayment



1
Expert's answer
2021-05-17T18:14:10-0400

The goods market is in equilibrium when aggregate demand is equal to the real output. The aggregate demand is the sum of consumption, investment, government expenditure and net exports. 

Y = AD 

Y = C+I + G + NX

Y = 100 + 0.75(Y-T) +200 + 150 + (200-(50+0.25Y))

Y = 100 + 0.75(Y-(50 + 0.5Y)) + 350 + 200 -50 -0.25Y

Y = 600 + 0.75(0.5Y -50) -0.25Y 

Y = 600 + 0.375Y -37.5 -0.25Y 

Y-0.125Y = 562.5 

0.875Y = 562.5 

    Y*=642.85


When exports decrease by 100 billion the real output decreases through the multiplier effect. 

Change in output = multiplier * change in exports 

            "= (\frac{1}{ 1-mpc}) \times100 \\

 = (\frac{1}{1-0.75}) \times 100\\

 = (\frac{1}{0.25}) \times 100\\

 = 400"

New real output (Y') = 642.85 - 400 

                 = 242.85

Before the change in exports:

Exports = 200 

Imports = 50+0.25Y

       = 50 + 0.25(642.85)

       = 50 + 160.7

       = 210.7

Net exports = exports - imports 

          = 200 - 210.7 

          = -10.7 


After the change in exports:

Exports = 100 

Imports = 50 + 0.25(242.85)

       = 110.7

Net exports = 100 - 110.7 

          = -10.7

As there is no change in net exports, there is no change in the current account hence balance of payment does not affect.